# The Map Equation Goes Neural: Mapping Network Flows with Graph Neural Networks

### v3

@@ C.4 Tabulated number of detected communities on real-world networks
Here, we report the average number of detected communities for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks. Table 4 lists the results on the Cora, CiteSeer, PubMed, Amazon Computer (PC), and Amazon Photo datasets. Table 5 lists the results on the Coauthor CS, Coatuhor Physics, Cora ML, Wiki CS, and OGB arxiv datasets.
@@
Table 8: Average number of detected communities and their standard deviations on real-world networks. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning.
@@
Method Arch. Cora CiteSeer PubMed PC Photo
Neuromap LIN $4.0\pm 1.2$ $4.5\pm 0.9$ $1.1\pm 0.3$ $1.0\pm 0.2$ $1.0\pm 0.0$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $3.7\pm 0.9$ $4.1\pm 1.1$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $9.7\pm 0.5$ $8.0\pm 0.2$
GIN $5.5\pm 0.8$ $5.1\pm 0.6$ $2.8\pm 0.4$ $3.6\pm 0.5$ $4.8\pm 0.9$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $6.5\pm 0.9$ $6.1\pm 0.5$
DMoN LIN $6.8\pm 0.6$ $5.8\pm 0.5$ $1.8\pm 0.8$ $4.0\pm 1.9$ $4.3\pm 1.4$
MLP $7.0\pm 0.2$ $6.0\pm 0.0$ $3.0\pm 0.0$ $9.8\pm 0.4$ $7.9\pm 0.3$
GCN $7.0\pm 0.0$ $6.0\pm 0.2$ $3.0\pm 0.2$ $10.0\pm 0.0$ $8.0\pm 0.0$
GIN $7.0\pm 0.0$ $5.9\pm 0.4$ $2.7\pm 0.5$ $10.0\pm 0.0$ $8.0\pm 0.0$
SAGE $7.0\pm 0.0$ $6.0\pm 0.2$ $3.0\pm 0.2$ $10.0\pm 0.0$ $8.0\pm 0.0$
NOCD LIN $7.0\pm 0.0$ $5.8\pm 0.4$ $1.7\pm 0.5$ $3.2\pm 1.1$ $2.8\pm 1.2$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $9.4\pm 0.9$ $7.4\pm 0.6$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$
GIN $6.2\pm 0.6$ $5.5\pm 0.6$ $2.9\pm 0.3$ $8.5\pm 0.6$ $6.8\pm 1.0$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.2$ $8.0\pm 0.0$
DiffPool LIN $2.2\pm 0.7$ $3.2\pm 0.6$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $2.2\pm 0.4$ $2.6\pm 0.7$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $5.3\pm 1.4$ $5.0\pm 0.7$
GIN $3.7\pm 0.6$ $3.0\pm 0.7$ $2.2\pm 0.4$ $4.6\pm 1.2$ $4.3\pm 1.1$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $3.2\pm 0.4$ $3.3\pm 0.5$
MinCut LIN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $1.3\pm 0.7$ $1.1\pm 0.3$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $7.7\pm 3.4$ $8.0\pm 0.0$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$
GIN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $6.1\pm 1.8$ $7.9\pm 0.3$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$
Ortho LIN $7.0\pm 0.2$ $5.6\pm 0.5$ $2.7\pm 0.4$ $4.5\pm 1.5$ $3.8\pm 1.4$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $8.3\pm 1.5$ $7.4\pm 0.7$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$
GIN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$
Infomap $282.2\pm 4.1$ $628.9\pm 2.7$ $924.9\pm 9.1$ $455.1\pm 5.1$ $223.5\pm 4.4$
Table 9: Average number of detected communities and their standard deviations on real-world networks. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning.

### v4

@@ D.6 Tabulated number of detected communities on real-world networks
Here, we report the average number of detected communities for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks for setting the number of hidden dimensions to $512$ and the maximum number of communities to $s=\left|Y\right|$.
@@
Table 9: Average number of detected communities and their standard deviations on real-world networks. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning. OOM stands for “out of memory”.
@@
Method Arch. Cora CiteSeer PubMed PC Photo CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $4.0\pm 1.2$ $4.5\pm 0.9$ $1.1\pm 0.3$ $1.0\pm 0.2$ $1.0\pm 0.0$ $10.6\pm 1.4$ $4.1\pm 0.7$ $2.6\pm 0.8$ $3.0\pm 0.9$ $1.2\pm 0.4$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $3.7\pm 0.9$ $4.1\pm 1.1$ $12.2\pm 1.2$ $5.0\pm 0.0$ $7.0\pm 0.0$ $7.6\pm 1.1$ $12.0\pm 2.4$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $9.7\pm 0.5$ $8.0\pm 0.2$ $13.8\pm 1.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $8.7\pm 0.9$ $16.2\pm 2.3$
GIN $5.5\pm 0.8$ $5.1\pm 0.6$ $2.8\pm 0.4$ $3.6\pm 0.5$ $4.8\pm 0.9$ $7.9\pm 0.8$ $3.9\pm 0.7$ $5.6\pm 0.8$ $4.9\pm 1.2$ $4.8\pm 1.5$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $6.5\pm 0.9$ $6.1\pm 0.5$ $12.7\pm 1.2$ $5.0\pm 0.0$ $7.0\pm 0.0$ $9.4\pm 0.6$ $13.0\pm 1.1$
DMoN LIN $5.9\pm 1.8$ $5.2\pm 1.5$ $3.0\pm 0.0$ $2.2\pm 0.9$ $2.3\pm 1.0$ $14.1\pm 2.5$ $4.5\pm 0.9$ $6.0\pm 1.8$ $5.1\pm 3.1$ $3.5\pm 3.2$
MLP $6.9\pm 0.3$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.2$ $10.0\pm 0.0$ $40.0\pm 0.0$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.2$ $7.0\pm 0.2$ $10.0\pm 0.0$ $40.0\pm 0.0$
GIN $7.0\pm 0.0$ $5.9\pm 0.3$ $3.0\pm 0.0$ $9.9\pm 0.3$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.2$ $39.5\pm 1.0$
SAGE $7.0\pm 0.2$ $5.9\pm 0.3$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $40.0\pm 0.0$
NOCD LIN $7.0\pm 0.0$ $5.8\pm 0.4$ $1.7\pm 0.5$ $3.2\pm 1.1$ $2.8\pm 1.2$ $10.8\pm 2.5$ $5.0\pm 0.0$ $7.0\pm 0.0$ $9.5\pm 0.7$ $10.1\pm 1.3$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $9.4\pm 0.9$ $7.4\pm 0.6$ $13.5\pm 1.1$ $5.0\pm 0.0$ $7.0\pm 0.2$ $9.8\pm 0.6$ $23.9\pm 3.6$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $36.2\pm 1.4$
GIN $6.2\pm 0.6$ $5.5\pm 0.6$ $2.9\pm 0.3$ $8.5\pm 0.6$ $6.8\pm 1.0$ $12.7\pm 1.4$ $4.8\pm 0.4$ $6.4\pm 0.6$ $9.4\pm 0.6$ $20.5\pm 3.2$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.2$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $31.8\pm 1.8$
DiffPool LIN $2.2\pm 0.7$ $3.2\pm 0.6$ $1.0\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$ $6.9\pm 2.5$ $3.5\pm 1.0$ $1.0\pm 0.0$ $2.5\pm 0.6$ OOM
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $2.2\pm 0.4$ $2.6\pm 0.7$ $11.6\pm 1.9$ $5.0\pm 0.0$ $7.0\pm 0.0$ $6.2\pm 0.9$ OOM
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $5.3\pm 1.4$ $5.0\pm 0.7$ $10.0\pm 1.2$ $5.0\pm 0.0$ $7.0\pm 0.0$ $5.5\pm 1.0$ OOM
GIN $3.7\pm 0.6$ $3.0\pm 0.7$ $2.2\pm 0.4$ $4.6\pm 1.2$ $4.3\pm 1.1$ $6.2\pm 3.1$ $4.2\pm 0.9$ $3.3\pm 0.9$ $3.9\pm 1.3$ OOM
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $3.2\pm 0.4$ $3.3\pm 0.5$ $8.4\pm 1.3$ $5.0\pm 0.0$ $7.0\pm 0.0$ $7.0\pm 0.8$ OOM
MinCut LIN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $1.3\pm 0.7$ $1.1\pm 0.3$ $13.8\pm 1.1$ $5.0\pm 0.0$ $7.0\pm 0.0$ $6.0\pm 2.4$ $10.8\pm 9.1$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $7.7\pm 3.4$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $36.2\pm 4.6$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $37.1\pm 5.4$
GIN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $6.1\pm 1.8$ $7.9\pm 0.3$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $8.4\pm 3.4$ $8.9\pm 3.8$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $40.0\pm 0.0$
Ortho LIN $7.0\pm 0.2$ $5.6\pm 0.5$ $2.7\pm 0.4$ $4.5\pm 1.5$ $3.8\pm 1.4$ $13.3\pm 0.6$ $5.0\pm 0.0$ $6.1\pm 1.4$ $7.8\pm 1.7$ $15.0\pm 1.7$
MLP $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $8.3\pm 1.5$ $7.4\pm 0.7$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $35.6\pm 1.8$
GCN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $40.0\pm 0.0$
GIN $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $40.0\pm 0.0$
SAGE $7.0\pm 0.0$ $6.0\pm 0.0$ $3.0\pm 0.0$ $10.0\pm 0.0$ $8.0\pm 0.0$ $15.0\pm 0.0$ $5.0\pm 0.0$ $7.0\pm 0.0$ $10.0\pm 0.0$ $39.9\pm 0.3$
Infomap $282.2\pm 4.1$ $628.9\pm 2.7$ $924.9\pm 9.1$ $455.1\pm 5.1$ $223.5\pm 4.4$ $834.4\pm 9.4$ $1160.8\pm 11.4$ $287.4\pm 3.8$ $747.0\pm 8.4$ $4840.0\pm 22.4$
Appendix E Performance impact of hidden layers and max. number of communities
Here, we report the performance difference in terms of achieved average AMI score between using (i) $4\sqrt{n}$ hidden layers and a maximum of $s=\sqrt{n}$ communities, and (ii) $512$ hidden layers and a maximum of $s=\left|Y\right|$ communities, that is, the “ground-truth” number of communities. The tabulated values show much better or worse the methods perform when using setup (ii) as compared to setup (i). That is, a positive value, shown in green, means that setup (ii) gives a better result while a negative value, shown in red, means that setup (i) gives a better result.
We find that, for Neuromap, NOCD, and DiffPool, setup (i) generally works better. In contrast, DMoN, MinCut, and Ortho often perform better with setup (ii), indicating that they require knowing the correct number of communities to perform well.
Table 10: Performance difference in terms of achieved average AMI score between (i) $4\sqrt{n}$ hidden layers and a maximum of $s=\sqrt{n}$ communities, and (ii) $512$ hidden layers and a maximum of $s=\left|Y\right|$ communities, that is, the “ground-truth” number of communities. OOM stands for “out of memory”.